OpenStudy (anonymous):

A 30-year-old worker plans to retire at age 65. He believes that $500,000 is needed to retire comfortably. How much should be deposited now at 7% compounded monthly to meet the $500,000 retirement goal? (Do not round until the final answer. Then, round to the nearest cent.)

OpenStudy (anonymous):

43455.01?

OpenStudy (anonymous):

Alright so similar to the last one, we need to solve for PV: $500,000=PV*(1+(.07/12))^(12*(65-30)) $500,000=PV*11.506 PV=$43,455.01

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

\[P(1+\frac{.07}{12})^{12\times 35}=500,000\] \[P=\frac{500,000}{(1+\frac{.07}{12})^{12\times 35}}\]